The consumption function is C=40+(0.8)Y. Let G=20 and NX=0. By how much does equilibrium output (Y^e) decline when in investments (I) falls from 20 to 0?
Answer #1

If I =20

AD= C+G+I+NX

Put AD= Y for finding equilibrium

Y= 40+0.8Y + 20+20

0.2Y= 80

Y = 400.

Now if I falls to 0,

Y = 40+0.8Y + 20

0.2Y = 60

Y = 300

So, equilibrium output falls by 400-300 = 100